A goldsmith has two alloys that are different purities of gold. The first is three-fourths pure gold and the second is five-twelfths pure gold. How many ounces of each should be melted and mixed in order to obtain a 6-oz mixture that is two-thirds pure gold?

Answer by checkley77(12844)

3/4x+5/12(6-x)=2/3*6
3x/4+30/12-5x/12=12/3
3x/4+2.5-5x/12=4
3x/4-5x/12=4-2.5
(3x*3-5x)/12=1.5
(9x-5x)/12=1.5
4x/12=1.5
x/3=1.5
x=4.5 oz. of the 3/4 pure gold is needed.
6-4.5=1.5 oz. of the 5/12 pure gold is needed.
Proof:
3/4*4.5+5/12*1.5=2/3*6
13.5/4+7.5/12=4
3.375+.625=4
4=4